OpenStudy (anonymous):

Convert the mixed number to an improper fraction. 2 and 7/10 A.14/10 B.27/10 C.34/10 D.37/10 Please Help!

OpenStudy (mathstudent55):

Are you able to see the mixed numeral below? \( 2 \dfrac{7}{10}\)

OpenStudy (mathstudent55):

\( 2 \dfrac{7}{10}\) \(= \dfrac{2 \times 10 + 7}{10} \)

OpenStudy (mathstudent55):

2 7/10 = 2 + 7/10 = 2/1 + 7/10 = (2 * 10)/(1 *10) + 7/10 = 20/10 + 7/10 = 27/10

OpenStudy (anonymous):

=27/10